OpenStudy (anonymous):

Find the maximum value of f(x) = -3x^2 + 12x-8

OpenStudy (anonymous):

\[f \prime (x)=-6x+12 \]To get the maximum, the 1st derivative should be equal to zero where the tangent is zero. \[0=-6x+12\]\[-6x=-12\]\[\frac{ -6x }{ -6 }=\frac{ -12 }{ -6 }\]\[x=2\]

OpenStudy (anonymous):

Using calculus, as suggested by lynn, or if you just graph the parabola,, you can find the vertex by letting x=-b/2a and substituting that value in to find what y is.

OpenStudy (anonymous):

@Lynncake , wasn't trying to correct you, just giving warsop an option in case he isn't in calculus or learned derivatives yet.

OpenStudy (anonymous):

it's fine :) this is an open study